# Visual Data Diagnosis and Debiasing with Concept Graphs

### v1

@@ Abstract
The widespread success of deep learning models today is owed to the curation of extensive datasets significant in size and complexity. However, such models frequently pick up inherent biases in the data during the training process, leading to unreliable predictions. Diagnosing and debiasing datasets is thus a necessity to ensure reliable model performance. In this paper, we present ConBias, a novel framework for diagnosing and mitigating Concept co-occurrence Biases in visual datasets. ConBias represents visual datasets as knowledge graphs of concepts, enabling meticulous analysis of spurious concept co-occurrences to uncover concept imbalances across the whole dataset. Moreover, we show that by employing a novel clique-based concept balancing strategy, we can mitigate these imbalances, leading to enhanced performance on downstream tasks. Extensive experiments show that data augmentation based on a balanced concept distribution augmented by ConBias improves generalization performance across multiple datasets compared to state-of-the-art methods. We will make our code and data publicly available.
@@ 4.1 Setup
Datasets. We use three datasets in our work: Waterbirds [45], UrbanCars [22], and COCO-GB [50], that are commonly used in the bias mitigation domain. We tackle background bias in the Waterbirds dataset, background and co-occuring object bias in the UrbanCars dataset, and finally gender bias in COCO-GB. All the tasks are binary classification tasks. More details on the training splits and class labels are provided in the supplementary material.

### v2

@@ Abstract
The widespread success of deep learning models today is owed to the curation of extensive datasets significant in size and complexity. However, such models frequently pick up inherent biases in the data during the training process, leading to unreliable predictions. Diagnosing and debiasing datasets is thus a necessity to ensure reliable model performance. In this paper, we present ConBias, a novel framework for diagnosing and mitigating Concept co-occurrence Biases in visual datasets. ConBias represents visual datasets as knowledge graphs of concepts, enabling meticulous analysis of spurious concept co-occurrences to uncover concept imbalances across the whole dataset. Moreover, we show that by employing a novel clique-based concept balancing strategy, we can mitigate these imbalances, leading to enhanced performance on downstream tasks. Extensive experiments show that data augmentation based on a balanced concept distribution augmented by ConBias improves generalization performance across multiple datasets compared to state-of-the-art methods. [^1] Code: https://github.com/rwchakra/conbias
@@ 3.4 Concept Graph Debiasing
Note on concept set annotations. We assume the availability of reliable ground truth concept sets. Such annotations already exist for the datasets we investigate - Waterbirds, UrbanCars, and COCO-GB. We agree that unreliable ground truth concept sets would hinder generalization abilities, but this assumption is not dissimilar to the assumption of reliable ground truth labels in classification tasks. Moreover, the reliance on ground truth concept sets, sometimes referred to as concept banks, have also been considered in prior work [57]. Ground truth concept sets serve as auxiliary knowledge bases and provide human level interpretability to the task at hand.
Note on computational complexity. In general, given $K$ classes and $C$ concepts, the graph clique enumeration is expected to grow in $O(exp(K+C))$. However, in practice, we find that constraining clique sizes $\leq 4$ leads to interpretable bias combinations, with no significant effect of the exponential runtime.
@@ 4.1 Setup
Datasets. We use three datasets in our work: Waterbirds [45], UrbanCars [22], and COCO-GB [50], that are commonly used in the bias mitigation domain. We tackle background bias in the Waterbirds dataset, background and co-occuring object bias in the UrbanCars dataset, and finally gender bias in COCO-GB. All the tasks are binary classification tasks. More details on the training splits and class labels are provided in the supplementary material. For Waterbirds, there are 66 nodes and 865 edges in the concept graph. For UrbanCars, the graph contains 19 nodes and 106 edges. For COCO-GB, there are 81 nodes and 2326 edges in the graph.
@@ 5 Conclusion, Limitations, and Future Works
Acknowledgements
This research is partially supported by a grant from Apple Inc. We thank Nicholas Apostoloff, Oncel Tuzel, and Jerremy Holland for their valuable feedback on the draft of this paper. Any views, opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and should not be interpreted as reflecting the views, policies or position, either expressed or implied, of Apple Inc. The authors would also like to thank the Norwegian Research Council for funding a doctoral research visit to the Human Sensing Lab, Carnegie Mellon University.
@@ Acknowledgements
  • Section 14: The sampling algorithm for attribute rebalancing.
  • Section 15: An illustration of our method diagnosing spurious correlations in ImageNet-1k.
@@ 14 Sampling Algorithm
The rebalance sampling algorithm 1 receives the concept graph as a concept co-occurrence matrix. The algorithm iterates through all cliques with an order of decreasing clique sizes to make sure we would not double compensate for the imbalance, e.g., 3-cliques would impact the already-balanced 2-cliques if we operate in a bottom-up fashion. For each iteration, it retrieves all cliques of concepts of size k along with their corresponding frequencies with each class. The algorithm identifies the maximum co-occurrence count among all classes for each combination and checks if any class is under-represented by comparing its count with the maximum. If a class is under-represented, the algorithm computes the number of synthetic samples needed to balance the representation and adds this information to the results list. This process continues for all combinations and classes until all clique sizes have been processed. The output of the algorithm is a list of queries specifying the class, concept combination, and the number of samples needed to balance the dataset.
@@
15 Diagnosing ImageNet-1k
In this section we demonstrate the usefulness of ConBias in diagnosing spurious concepts in a more complex dataset, i.e. ImageNet-1k. Specifically, we investigate the classes ambulance, beach wagon, sports car, limousine, minivan, jeep, convertible, cab. These are associated with the superclass of Car. This approach can be used for any set of classes in the dataset. We use the open-source concept annotations available here.
(a) Imbalance for 1-clique Tram
(b) Imbalance for 1-clique Prison
(c) Imbalance for 2-clique (Golf, Sorrel)
(d) Imbalance for 2-clique (Ski, Grille)
Figure 17: Concept imbalances for cars in ImageNet-1k
In Figure 17 we notice some interesting imbalances discovered by ConBias. For 1-clique concept combinations such as Tram and Prison, the dataset disproportionately contains images of cabs and ambulances respectively. For 2-clique concept combinations, the dataset disproportionately represents the jeep, convertible, and beach wagon classes. It is evident that such concepts are spurious when it comes to classifying a car type, but a strong imbalanced distribution would bias the classifier to pick up on spurious features. In this way, ConBias helps us uncover such biases, allowing for intervention in downstream tasks.